OpenStudy (anonymous):

answer for 4x-y=2 and 5x-2y=-2

OpenStudy (saifoo.khan):

simultanious?

OpenStudy (anonymous):

through substitution

OpenStudy (saifoo.khan):

x = 2 y = 6

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

Welks.

OpenStudy (saifoo.khan):

Please hit good answer.

OpenStudy (saifoo.khan):

Thnks! :)

OpenStudy (anonymous):

4x-y=2 and 5x-2y=-2 trough substitution

OpenStudy (anonymous):

4x-y=2 5x-2y=-2 4x-2=y so.. 5x-2(4x-2)=-2 5x-8x+4=-2 5x-8x=-2-4 -3x=-6 x=2 :) then... find y through substitution if x=2 4(2)-y=2 8-y=2 y=6 :)